OpenStudy (anonymous):

How to convert x=3+3cos(theta) and y=2+5sin(theta) to rectangular by eliminating the parameter

OpenStudy (anonymous):

\[ \frac{x-1}3=\cos(\theta) \]

OpenStudy (anonymous):

\[ \sin(\theta) =\pm\sqrt{1-\cos^2(\theta)} = \pm \sqrt{1-\left(\frac{x-1}3\right)^2} \]

OpenStudy (anonymous):

\[ y=2\pm 5\sqrt{1-\left(\frac{x-1}3\right)^2} \]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

This parametrization can only be regularized for certain intervals of \(\theta\), otherwise it stops being a function.

OpenStudy (anonymous):

There may be a better way of doing this..

OpenStudy (anonymous):

\[ \frac{x-1}{3}=\cos(\theta)\\ \frac{y-2}{5} = \sin(\theta)\\ \left(\frac{x-1}{3}\right)^2+\left(\frac{y-2}{5}\right)^2 = \cos^2(\theta)+\sin^2(\theta) = 1 \]

OpenStudy (anonymous):

So a less pelletty answer is just\[ \left(\frac{x-1}{3}\right)^2+\left(\frac{y-2}{5}\right)^2 = 1 \]

OpenStudy (anonymous):

Medal now.